Does the sequence 3+ (-1)^n converge to 2. Justify

1
Expert's answer
2022-02-09T11:54:12-0500

The sequence "x_n=3+(-1)^n" does not converge to 2 because its infinite subsequence "x_{2n}=3+(-1)^{2n}=3+1=4" does not converge to 2.